Alecs is 6 years older than Sistine. In 2 years, Alecs will be twice as old as Sistine. Find their ages
Answer by JulietG(1812)

A = S + 6 [Alecs is 6 years older than Sistine.]
A + 2 = 2S [In 2 years, Alecs will be twice as old as Sistine.]
.
Substitute the value of A in the first equation for A in the second.
(S+6) + 2 = 2S
Add
S + 8 = 2S
If 8 and S = 2 * S, then S is 8.
Is S is 8, then A is 14.